Genes A and B are located 10cM from each other on a chromosome. Gene C is located 25cM from gene A and 15cM from gene B. Assuming that I = 0, what is the probability that the trihybrid ABC/abc will produce an ABC gamete?

Genes A and B are located 10cM from each other on a chromosome. Gene C is located 25cM from gene A and 15cM from gene B. Assuming that I = 0, what is the probability that the trihybrid ABC/abc will produce an ABC gamete?

38.25%
25%
1.5%
76.5%

.75%

Please include explanation!
Answer #1

Distance between genes A and B = 15 cM

Distance between genes B and C = 10 cM

The probability recombination

Probability of SCO - DCO = 0.25 - 0.015 = 0.235

Probability of non- recombinants = 1 - 0.235 = 0.765

Half of the non recombinants are ABC and the other half are abc. Therefore, the probability that an ABC gamete is produced is 0.5 X 0.765 = 0.3825 or 38.25%

Therefore, the probability of producing ABC gamete is 38.25%